Kelso Electric is an all-equity firm with 54,500 shares of stock outstanding. The company is considering the issue of $370,000 in debt at an interest rate of 8 percent and using the proceeds to repurchase stock. Under the new capital structure, there would be 34,000 shares of stock outstanding. Ignore taxes. What is the break-even EBIT between the two plans?
Answer #1

at break even EBIT , EPS of both the options would be equal.

let x be the break even EBIT,

at this point,

=> x / 54,500 shares = (x- 29,600) / 34,000 shares................(interest on debt = 370,000*8%=>29,600).

=.34,000 x = 54,500 x - 1,613,200,000

=>1,613,200,000 = 20,500x

=>x =78,693....(rounded to nearest dollar).

Break even EBIT = 78,693.